OpenStudy (anonymous):

medal & fan Which shows the numbers in order from least to greatest? 7.8 , 7.6 , 7.41 A. 7.8 , 7/41 ,7.6 B . 7.8 , 7.6 , 7.41 C . 7.41 , 7.6 , 7.8 D. 7.41 , 7.8 , 7.6

OpenStudy (anonymous):

Basically just take out the 7 and then imagine them as if they weren't decimals but actual integers. In other words, order the numbers 80,60, and 41 in order from least to greatest.

OpenStudy (anonymous):

yeah but what is it i dont get i

OpenStudy (anonymous):

You take the smallest number and put it first and then order the numbers so that smaller numbers are first and larger numbers are last. Which of 80, 60, and 41 is the smallest?

OpenStudy (anonymous):

so it is ...

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@LegendarySadist i need the anwser for u to get the medal &fan

OpenStudy (anonymous):

Like I said, arrange 80, 60, and 41 in order of smallest to largest to find the answer.

OpenStudy (anonymous):

so 40 60 80

OpenStudy (anonymous):

so it would be c

OpenStudy (anonymous):

Yes, it would be c :)

OpenStudy (anonymous):

can u help me with one more

OpenStudy (anonymous):

Sure, I can help with one more

OpenStudy (anonymous):

Which mixed number is greater than 8.45?

OpenStudy (anonymous):

A. 8 3/8 B. 8 4/9 C. 8 3/5 D. 8 5/12

OpenStudy (anonymous):

Well I'll turn them into decimal form and you can go from there 8 3/8 = 8.37 8 4/9 = 8.44 8 3/5 = 8.60 8 5/12 = 8.41 Which number after the 8 is bigger than 45?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

so 8 3/5

OpenStudy (anonymous):

Yep, that would be it. Good job :)

OpenStudy (anonymous):

tanks your the best :)

OpenStudy (anonymous):

haha, no problem :) have a nice day

OpenStudy (anonymous):

u too and tank you again